HW-Practice-Questions In Exercises Questions 82, 90, and 92, begin by graphing the absolute value function, f(x) = Then use transformations of this graph to graph the given function 82.) g(x) -X +3 90.) g(x)x+42 92.) g(x 2x+3
Answer #1

82. g(x) = \left | x \right |+3

For getting graph of g(x) we shifted graph of f(x) = \left | x \right | by 3 unit upward direction.

90. g(x) = -\left | x+4 \right |+2

Firstly we shifted graph of f(x) = \left | x \right | by 4 unit in left direction. Then we inverted it about x axis. Then shift the graph by 2 unit in upward direction.